7. An oxygen tank reads 130 kPa with a temperature of 25 °C. What will the pressure read if the tank is cooled to -183 °C?
Answer #1

Given:

Pi = 130 kPa

Ti = 25.0 oC

= (25.0+273) K

= 298 K

Tf = -183.0 oC

= (-183.0+273) K

= 90 K

use:

Pi/Ti = Pf/Tf

130 kPa / 298.0 K = Pf / 90.0 K

Pf = 39.2617 kPa

Answer: 39.3 kPa